Does there pass a plane through the lines (x+4)÷3 =y÷2 =(z−1)÷3 and x÷2 =(y−1)÷1 =(z+1)÷1?
Justify.
1
Expert's answer
2020-06-15T07:03:23-0400

Solution

The plane passes through the lines if the lines are parallel or intersect. The line "(x+4)\div3=y\div2=(z-1)\div3" has a direction vector"\overrightarrow{s}" (3, 2, 3) and a point, that belongs to it, is A(-4, 0, 1).

The line "x\div2=(y-1)\div1=(z+1)\div1" has a direction vector "\overrightarrow{n}" (2, 1, 1) and a point, that belongs to it, is B(0, 1, -1).

If the lines are parallel, then the coordinates of their direction vectors are proportional:

"3\div2\not =2\div1\not =3\div1" .

the lines are not parallel.

If the lines intersect, the vectors "\overrightarrow{s}", "\overrightarrow{n}", and "\overrightarrow{AB}" are complanar,

"\overrightarrow{s}*\overrightarrow{n}*\overrightarrow{AB}=0" ,

"\begin{vmatrix}
 0+4 & 1-0&-1-1 \\
 3 & 2&3\\
2&1&1
\end{vmatrix}=\begin{vmatrix}
 4 & 1&-2 \\
 3 & 2&3\\
2&1&1
\end{vmatrix}=\\=4*2*1+3*1*(-2)+1*3*2-(-2)*2*2-1*3*1-3*1*4=\\=8-6+6+8-3-12=1\not =0"

The lines do not intersect.

Answer:

the plane does not pass through the lines.